OpenStudy (anonymous):

For $360, a rock-climbing gym offers a yearly membership where members can climb as many days as they want and pay $4 per day for equipment rental. Nonmembers pay $10 per day to use the gym and $6 per day for equipment rental. Find the total number of visits when the cost for a member and nonmember will be the same

OpenStudy (anonymous):

@rajathsbhat

OpenStudy (anonymous):

ok let's model an equation which will tell us how much you'll be shelling out if you're a) a member and b) a non member. Suppose you're a member. You first give $360 and $4 for every day (for the eqpt) If you only go there for one day in that year, you'll be paying 360 + 4 If you go there for 2 days, you pay 360 + 4 + 4 For 5 days, you pay 360 + 4 + 4+ 4 + 4+ 4 = 360 + 4(5). Now, if you go there for 'x' days in that year, you pay 360 + 4(x). Do you get it?

OpenStudy (anonymous):

yeah ido so ill do 360(4) then add it to 360

OpenStudy (anonymous):

no no you can do that if you go there for the whole year. But you're not doing that.

OpenStudy (anonymous):

the member will pay 360 for the "membership" PLUS 4 dollars for everyday. The non-member will pay $16 everyday. So you want to see how many days it'll be until a member and non-member end up paying the same amount. You know what i mean?

OpenStudy (anonymous):

ohhh i see

OpenStudy (anonymous):

A. 480 days B. 175 days C. 60 days D. 30 days

OpenStudy (anonymous):

yeah :) the non-member will pay only $16 if he goes there for only one day in that year. He'll pay $32 for 2 days i.e. 16(2) If he goes there for x days, he'll pay 16(x) okay? so to find the no of days till they pay the same thing you equate the two models. 360+4(x)=16(x) Finding x will mean finding the no of days...

OpenStudy (anonymous):

Tell me what you get for x. I'm interested.

OpenStudy (anonymous):

D 30 days thanks

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

ohh yeah i forgot wat is plottinglike in graphing

OpenStudy (anonymous):

plotting is just putting points on a graph

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

:)